When Molybdenum(VI) Oxide and Zinc are heated together they react to form 

molybdenum (III) Oxide and Zinc (II) Oxide. What mass of Molybdenum (III) 

Oxide and Zinc (II) Oxide is formed when 20.0 grams of Molybdenum(VI) Oxide is 

reacted with 10.0 grams of Zn?


1
Expert's answer
2022-05-03T14:00:01-0400

2MoO2 + Zn = Mo2O3 + ZnO


n(MoO2)=m(MoO2)/M(MoO2)=20/128=0.156 mol


n(Zn)=m(Zn)/M(Zn)=10/65=0.154 mol


n(MoO2) : n(Zn) = 2 : 1, so n(Zn) excess


n(MoO2) : n(Mo2O3) = 2 : 1


n(Mo2O3) = n(MoO2)/2 = 0.156/2 = 0.078 mol


n(MoO2) : n(ZnO) = 2 : 1


n(ZnO) = n(MoO2)/2 = 0.156/2 = 0.078 mol


m(Mo2O3) = 0.078×240 = 18.7 g


m(ZnO) = 0.078×81 = 6.3 g